Find the Slope, if it exist. 4X - 10Y = 20
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

Find the Slope, if it exist. 4X - 10Y = 20
----------------------
Put the equation into slope-intercept form:
10y = 4x - 20
y = (4/10)y - 2
----------------------
slope = 4/10 = 2/5
=============================
Cheers,
Stan H.

Answer by Alan3354(69443)

For standard from eqns, ax + by = c,
m = -a/b
--------------
m = -4/-10 = 2/5 or 0.4